# Multilevel neural simulation-based inference

### v1

@@ B.2 Proof of Theorem 2
where $C_{l}$ is the cost of evaluating $f^{l}$ at level $l$. In addition, we also have the following upper bound on the variance using Theorem 1:
@@
such $\quad\text{such that}\quad\text{Cost}({\ell}_{\text{MLMC}}(\phi);n_{0},\ldots,n_{L})\leq C_{\text{budget}},$
@@
$\quad\Leftrightarrow\quad n^{\star}_{l}=\sqrt{\frac{K_{l}(\phi)}{\nu(C_{l}+C_{l+1})}\left(\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1\right)}.$ (40)
@@
Plugging the results of 39 and 40 into 41, we get:
@@
$=\sqrt{\frac{K_{0}(\phi)}{\nu C_{0}}\left(\left\|G^{0}\right\|^{4}_{W^{1,4}}+1\right)}\times C_{0}+\sum_{l=1}^{L}\sqrt{\frac{K_{l}(\phi)}{\nu(C_{l}+C_{l+1})}\left(\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1\right)}\times(C_{l}+C_{l-1})$
$=\nu^{-\frac{1}{2}}\left(\sqrt{K_{0}(\phi)C_{0}\left(\left\|G^{0}\right\|_{W^{1,4}}^{4}+1\right)}+\sum_{l=1}^{L}\sqrt{K_{l}(\phi)(C_{l}+C_{l+1})\left(\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1\right)}\right),$
@@
$\nu=\frac{\left(\sqrt{K_{0}(\phi)C_{0}\left(\left\|G^{0}\right\|_{W^{1,4}}^{4}+1\right)}+\sum_{l=1}^{L}\sqrt{K_{l}(\phi)(C_{l}+C_{l+1})\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}+1\right)}\right)^{2}}{C_{\text{budget}}^{2}}.$ (42)
We can now use the expression for $\nu$ that we obtained in 42 to obtain a simplified expression for $n_{0},\ldots,n_{L}$ (using 39 and 40):
$n_{0}^{\star}$ $\propto\frac{C_{\text{budget}}}{\sqrt{C_{0}}}\sqrt{\left\|G^{0}\right\|_{W^{1,4}}^{4}+1},\quad\quad n_{l}^{\star}$ $\propto\frac{C_{\text{budget}}}{\sqrt{C_{l}+C_{l+1}}}\sqrt{\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1}\quad\text{for }l\in\{1,\ldots,L\}.$
@@
Appendix C Experimental details & additional results
For all the experiments, we use a Mac M4 CPU with $16$ GB memory for the training of neural networks. Running all the experiments roughly takes half a day. For optimisation, we use batch gradient descent with the Adam optimiser (Kingma and Ba, 2015) using Pytorch (Paszke et al., 2017). For the construction of the conditional density estimator, we use the SBI package (Tejero-Cantero et al., 2020). We also use the BayesFlow package (Radev et al., 2023b) for some of the figures. For each experiment, we fix the number of epochs unless stated otherwise. We use the same setting and stopping criterion for the NPE and the NLE baseline as in the default implementation of the SBI package. We set the learning rate to $10^{-4}$ for all the experiments.
C.1 The g-and-k distribution
@@
where $\Phi^{-1}$ is the quantile function of the standard normal distribution and $\text{erf}^{-1}$ is the inverse of error function. We set the prior distribution to be a tensor product of marginal distributions on each parameter: $\theta_{1},\theta_{2},\theta_{3}\sim\text{Unif}([0,3]^{3})$ and $\theta_{4}\sim\text{Unif}([0,\exp(0.5)])$. Note that we always resort to an approximation method for the evaluation of $\text{erf}^{-1}(\cdot)$. For the high-fidelity simulator, we use an accurate approximation implemented in Scipy (Virtanen et al., 2020). For the low-fidelity simulator, we use a Taylor expansion of $\text{erf}^{-1}$ up to the third order:
@@
An advantage of the g-and-k distribution is that its density can be approximated numerically almost exactly. Following Rayner and MacGillivray (2002), we first numerically obtain $G_{\theta}^{-1}(x)$ by solving numerically for $x_{i}-G_{\theta}(u_{i})=0$ using a root-solving algorithm (Press, 2007)111 Note that since $G_{\theta}(u)$ is a quantile function, $F_{\theta}(x):=G_{\theta}^{-1}(x)$ is a cumulative distribution function., then we obtain the density function by taking $\hat{p}(x\,|\,\theta):=F^{\prime}_{\theta}(x)=\partial F_{\theta}(x)/\partial x$. For this second step, we typically use a finite difference approximation. Seed-matching is simply done by using same the $u$ and $\theta$.
@@
We use the neural spline flow (NSF) (Durkan et al., 2019). We pick $10$ bins, span of $[-7,7]$ and $1$ coupling layer since we only have one dimensional input. The conditioner for the NSF is a multilayer perceptron neural network (MLP) with $3$ hidden layers of 50 units, and $10\%$ dropout, trained for $10,000$ epochs.
@@
We use NSF with $3$ bins, span of $[-3,3]$ and $3$ coupling layers. The conditioner for the NSF is a MLP with $2$ hidden layers of $50$ units, and $10\%$ dropout, trained for $800$ epochs. For each true parameter values, we produce $m=1000$ i.i.d samples and obtain quantile based $4$-dimensional summary statistics following Prangle (2016).
@@
We calculate the forward KL-divergence between the almost exact density and the approximated density over $2000$ equidistant points in $[-30,30]$.
@@
We calculate NLPD over $500$ simulations. Empirical coverage is estimated using average value of $500$ simulated datasets, where we draw $2000$ posterior samples from each simulations. We then calculate $1-\beta$-highest posterior density credible interval, where $\beta$ is $101$ equidistant points between 00 and $1$.
@@
Figure 8: Additional results for the g-and-k experiment. (a) Histogram of $1000$ seed-matched samples from low- and high-fidelity simulator. (b) ISE (integrated squared error) for NLE ($\downarrow$): sum of the squared distance between (almost) exact density and approximated density over $2000$ points in the interval $[-30,30]$. The performance of ML-NLE improves as $n_{1}$ increases. (c) Approximated density and (almost) exact density for NLE across four simulations. The first row of (c) shows that we get better approximation of the density when combining low and fidelity samples than using them separately. The second row of (c) shows the improvement of the performance of our ML-NLE as $n_{1}$ increases. These results demonstrate the effectiveness of our method. (d) Recovery plot for NPE: It measures how well the ground truth value is captured by the median of approximated posterior. The x-axis shows the ground truth parameter value and the y-axis shows the median of the posterior distribution with the median absolute deviation, i.e. ±$\pm\text{median}(|\theta-\text{median}(\theta)|)$, shown around the points. Here $r$ and $R^{2}$ denotes the correlation coefficient ($\uparrow$) and the coefficient of determination ($\uparrow$) between the ground truth values and the estimated median, respectively. The dashed diagonal line indicates perfect recovery of the true parameter. Using MLMC leads to significant improvements in the parameter recovery, which become more pronounced as $n_{1}$ increases.
C.2 Training time comparison
We report training time per epoch for our multilevel versions of NLE and NPE and their standard counterparts for the g-and-k experiment; see Table 1. For both methods, we picked $n=2000$ for the standard NLE/NPE with MC loss and $n_{0}=1000,n_{1}=500$ for our ML-NLE/ML-NPE with MLMC loss such that the total number of samples are the same. Each network is trained independently $10$ times to assess uncertainty, with a training budget of $500$ epochs for NLE and $100$ epochs for NPE.
Method Training Time per Epoch (s ×$\times 10^{-3}$)
@@ C.2 Training time comparison
C.3 Ornstein–Uhlenbeck process
@@
We use NSF with $2$ bins, span of $[-2,2]$ and $2$ coupling layers. The conditioner for the NSF is a MLP with $1$ hidden layer of $20$ units, and $20\%$ dropout. For ML-NPE, we trained the network for $500$ epochs. For TL-NPE, we set a high maximum number of epochs and used on early stopping. The stopping criterion is based on the validation loss computed on $10\%$ of the data, with a patience parameter controlling how many epochs to wait before stopping. For each true parameter values, we produce $m=1$ sample and use $5$ representative data points as a summary statistics. The subsamples are taken at equal interval in log space such that we take $0,3,10,31,99$th data points.
@@
Figure 9: Additional results for the OUP experiment. (a) Example of one seed-matched sample from high and low fidelity simulators. (b) Recovery plot for the reference NPE posterior with $n=10^{4}$. (c)-(f) Recovery plots for ML-NPE and TL-NPE for $n_{1}=10$ and $n_{1}=100$. For TL-NLE, we picked the result corresponding to the best performing hyperparameter of their stopping criterion (i.e. patience $=1$ for $n_{1}$ = 10 and patience $=20$ for $n_{1}=100$) to visualise. For $n_{1}=10$, ML-NPE achieves better performance on $\gamma$ and $\mu$. For $n_{1}=100$, it achieves superior performance on $\gamma$ and competitive results on the remaining two. Recall that, unlike TL-NPE, ML-NPE does not have any hyperparameter to tune.
C.4 Toggle switch model
@@
We follow the implementation by Key et al. (2025). Given parameters $\theta=[\alpha_{1},\alpha_{2},\beta_{1},\beta_{2},\mu,\sigma,\gamma]^{\top}$, we can sample from the simulator by $x\sim\widetilde{\mathcal{N}}(\mu+u_{T},\mu\sigma/u_{T}^{\gamma})$, where $\widetilde{\mathcal{N}}$ denotes the truncated normal distribution on $\mathbb{R}_{+}$. Here, $u_{T}$ is given by the discretised-time equation as follows:
@@
$C_{\text{ML-NLE}}=c\left(n_{0}T_{0}+\sum_{l=1}^{L}n_{l}(T_{l}+T_{l-1})\right)\quad\text{and}\quad C_{\text{NLE}}=cTn.$
This yields a total data generation cost for ML-NLE with $(n_{0},n_{1},n_{2})=(10^{4},500,300)$, $(T_{0},T_{1},T_{2})=(50,80,300)$ as $C_{\text{MLMC}}=60300$, assuming a unit cost $c=1$. We then allocate the same computational budget to NLE at different fidelities. This results in the following sample sizes: $n_{\text{high}}=C_{\text{ML-NLE}}/T_{2}=2010$, $n_{\text{med.}}=C_{\text{ML-NLE}}/T_{1}=7537$, and $n_{\text{low}}=C_{\text{ML-NLE}}/T_{0}=12060$.
Note that $\mathcal{U}$ varies with the fidelity level of the simulator, as it is given by $\mathbb{R}^{2T+1}$. This discrepancy can complicate seed matching across fidelities. To address this, we can define a unified domain for the noise $\mathcal{U}=\bigcup_{l=0}^{L}\mathcal{U}^{l}$ shared across all fidelity levels, and assume that for lower fidelities (smaller $l$), certain components of the input are simply disregarded. For seed matching, we share $\theta$ and $u_{l}\cap u_{l-1}=u_{l-1}$ where $u_{l}\in\mathcal{U}^{l}\subseteq\mathcal{U},\forall l$.
@@
We use Gaussian mixture density network (Bishop, 1994) with two mixture components. We use MLP with $2$ hidden layers and $20$ units to estimate the parameters: means, variances, and mixture weights, trained for $10,000$ epochs.
@@
Figure 10: Additional results for the Toggle-switch experiment. (a) Histograms of $1000$ seed-matched samples from the high-, the medium-, and the low-fidelity simulators. (b) Approximated density and samples from high-fidelity simulator across four simulations. ML-NLE demonstrates superior performance compared to the NLE baselines trained exclusively on high-, medium-, and low-fidelity data with same total simulation budget. In particular, the density estimated by ML-NLE aligns more closely with high-fidelity simulations, highlighting the improved accuracy of our method.
C.5 Cosmological simulations

### v2

@@ A.2 Proof of Theorem˜2
where $C_{l}$ is the cost of evaluating $f^{l}$ at level $l$. In addition, we also have the following upper bound on the variance using Theorem˜1:
@@
$\quad\text{such that}\quad\text{Cost}({\ell}_{\text{MLMC}}(\phi);n_{0},\ldots,n_{L})\leq C_{\text{budget}},$
@@
$\quad\Leftrightarrow\quad n^{\star}_{l}=\sqrt{\frac{K_{l}(\phi)}{\nu(C_{l}+C_{l-1})}\left(\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1\right)}.$ (41)
@@
Plugging the results of ˜40 and ˜41 into ˜42, we get:
@@
$=\sqrt{\frac{K_{0}(\phi)}{\nu C_{0}}\left(\left\|G^{0}\right\|^{4}_{W^{1,4}}+1\right)}\times C_{0}$
$\quad+\sum_{l=1}^{L}\sqrt{\frac{K_{l}(\phi)}{\nu(C_{l}+C_{l-1})}\left(\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1\right)}\times(C_{l}+C_{l-1})$
$=\nu^{-\frac{1}{2}}\left(\sqrt{K_{0}(\phi)C_{0}\left(\left\|G^{0}\right\|_{W^{1,4}}^{4}+1\right)}\right$.
$\quad\left.+\sum_{l=1}^{L}\sqrt{K_{l}(\phi)(C_{l}+C_{l-1})\left(\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1\right)}\right),$
@@
$\nu=\frac{\left(\sqrt{K_{0}(\phi)C_{0}\left(\left\|G^{0}\right\|_{W^{1,4}}^{4}+1\right)}+\sum_{l=1}^{L}\sqrt{K_{l}(\phi)(C_{l}+C_{l-1})\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}}+1\right)}\right)^{2}}{C_{\text{budget}}^{2}}.$ (43)
We can now use the expression for $\nu$ that we obtained in ˜43 to obtain a simplified expression for $n_{0},\ldots,n_{L}$ (using ˜40 and ˜41):
$n_{0}^{\star}$ $\propto\frac{C_{\text{budget}}}{\sqrt{C_{0}}}\sqrt{\left\|G^{0}\right\|_{W^{1,4}}^{4}+1},\quad\quad n_{l}^{\star}$ $\propto\frac{C_{\text{budget}}}{\sqrt{C_{l}+C_{l-1}}}\sqrt{\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{2}+1}\quad\text{for }l\in\{1,\ldots,L\}.$
@@
Appendix B Experimental details & additional results
For all the experiments, we use a Mac M4 CPU with $16$ GB memory for the training of neural networks. Running all the experiments roughly takes half a day. For optimisation, we use batch gradient descent with the Adam optimiser [Kingma and Ba, 2015] using Pytorch [Paszke et al., 2017]. For the construction of the conditional density estimator, we use the SBI package [Tejero-Cantero et al., 2020]. We also use the BayesFlow package [Radev et al., 2023b] for some of the figures. For each experiment, we fix the number of epochs unless stated otherwise. We use the same setting and stopping criterion for the NPE and the NLE baseline as in the default implementation of the SBI package. We set the learning rate to $10^{-4}$ for all the experiments.
B.1 The g-and-k distribution
@@
where $\Phi^{-1}$ is the quantile function of the standard normal distribution and $\text{erf}^{-1}$ is the inverse of error function. We set the prior distribution to be a tensor product of marginal distributions on each parameter: $\theta_{1},\theta_{2},\theta_{3}\sim\text{Unif}([0,3]^{3})$ and $\theta_{4}\sim\text{Unif}([0,\exp(0.5)])$. Note that we always resort to an approximation method for the evaluation of $\text{erf}^{-1}(\cdot)$. For the high-fidelity simulator, we use an accurate approximation implemented in Scipy [Virtanen et al., 2020]. For the low-fidelity simulator, we use a Taylor expansion of $\text{erf}^{-1}$ up to the third order:
@@
An advantage of the g-and-k distribution is that its density can be approximated numerically almost exactly. Following Rayner and MacGillivray [2002], we first numerically obtain $G_{\theta}^{-1}(x)$ by solving numerically for $x_{i}-G_{\theta}(u_{i})=0$ using a root-solving algorithm [Press, 2007]111 Note that since $G_{\theta}(u)$ is a quantile function, $F_{\theta}(x):=G_{\theta}^{-1}(x)$ is a cumulative distribution function., then we obtain the density function by taking $\hat{p}(x\,|\,\theta):=F^{\prime}_{\theta}(x)=\partial F_{\theta}(x)/\partial x$. For this second step, we typically use a finite difference approximation. Seed-matching is simply done by using same the $u$ and $\theta$.
@@
We use the neural spline flow (NSF) [Durkan et al., 2019]. We pick $10$ bins, span of [7,7][-7,7] and $1$ coupling layer since we only have one dimensional input. The conditioner for the NSF is a multilayer perceptron neural network (MLP) with $3$ hidden layers of 50 units, and $10\%$ dropout, trained for $10,000$ epochs.
@@
We use NSF with $3$ bins, span of [3,3][-3,3] and $3$ coupling layers. The conditioner for the NSF is a MLP with $2$ hidden layers of $50$ units, and $10\%$ dropout, trained for $800$ epochs. For each true parameter values, we produce $m=1000$ i.i.d samples and obtain quantile based $4$-dimensional summary statistics following Prangle [2016].
@@
We calculate the forward KL-divergence between the almost exact density and the approximated density over $2000$ equidistant points in [30,30][-30,30].
@@
We calculate NLPD over $500$ simulations. Empirical coverage is estimated using average value of $500$ simulated datasets, where we draw $2000$ posterior samples from each simulations. We then calculate $1-\beta$-highest posterior density credible interval, where $\beta$ is $101$ equidistant points between 0 and $1$.
@@
Figure 6: Additional results for the g-and-k experiment. (a) Histogram of $1000$ seed-matched samples from low- and high-fidelity simulator. (b) ISE (integrated squared error) for NLE ($\downarrow$): sum of the squared distance between (almost) exact density and approximated density over $2000$ points in the interval [30,30][-30,30]. The performance of ML-NLE improves as $n_{1}$ increases. (c) Approximated density and (almost) exact density for NLE across four simulations. The first row of (c) shows that we get better approximation of the density when combining low and fidelity samples than using them separately. The second row of (c) shows the improvement of the performance of our ML-NLE as $n_{1}$ increases. These results demonstrate the effectiveness of our method. (d) Recovery plot for NPE: It measures how well the ground truth value is captured by the median of approximated posterior. The x-axis shows the ground truth parameter value and the y-axis shows the median of the posterior distribution with the median absolute deviation, i.e. $\pm\text{median}(|\theta-\text{median}(\theta)|)$, shown around the points. Here $r$ and $R^{2}$ denotes the correlation coefficient ($\uparrow$) and the coefficient of determination ($\uparrow$) between the ground truth values and the estimated median, respectively. The dashed diagonal line indicates perfect recovery of the true parameter. Using MLMC leads to significant improvements in the parameter recovery, which become more pronounced as $n_{1}$ increases.
B.2 Training time comparison
We report training time per epoch for our multilevel versions of NLE and NPE and their standard counterparts for the g-and-k experiment; see Table˜1. For both methods, we picked $n=2000$ for the standard NLE/NPE with MC loss and $n_{0}=1000,n_{1}=500$ for our ML-NLE/ML-NPE with MLMC loss such that the total number of samples are the same. Each network is trained independently $10$ times to assess uncertainty, with a training budget of $500$ epochs for NLE and $100$ epochs for NPE.
Method Training Time per Epoch (s $\times 10^{-3}$)
@@ B.2 Training time comparison
B.3 Ornstein–Uhlenbeck process
@@
We use NSF with $2$ bins, span of [2,2][-2,2] and $2$ coupling layers. The conditioner for the NSF is a MLP with $1$ hidden layer of $20$ units, and $20\%$ dropout. For ML-NPE, we trained the network for $500$ epochs. For TL-NPE, we set a high maximum number of epochs and used on early stopping. The stopping criterion is based on the validation loss computed on $10\%$ of the data, with a patience parameter controlling how many epochs to wait before stopping. For each true parameter values, we produce $m=1$ sample and use $5$ representative data points as a summary statistics. The subsamples are taken at equal interval in log space such that we take $0,3,10,31,99$th data points.
@@
Figure 7: Additional results for the OUP experiment. (a) Example of one seed-matched sample from high and low fidelity simulators. (b) Recovery plot for the reference NPE posterior with $n=10^{4}$. (c)-(f) Recovery plots for ML-NPE and TL-NPE for $n_{1}=10$ and $n_{1}=100$. For TL-NLE, we picked the result corresponding to the best performing hyperparameter of their stopping criterion (i.e. patience =1=1 for $n_{1}$ = 10 and patience =20=20 for $n_{1}=100$) to visualise. For $n_{1}=10$, ML-NPE achieves better performance on $\gamma$ and $\mu$. For $n_{1}=100$, it achieves superior performance on $\gamma$ and competitive results on the remaining two. Recall that, unlike TL-NPE, ML-NPE does not have any hyperparameter to tune.
B.3.1 Experiment with different dimension of $\theta$
We now conduct an experiment to explore a failure mode of our method. We include an additional parameter in the high fidelity simulator, termed “initial value” $\theta_{4}=x_{0}\sim\text{uniform}([0,4])$, instead of fixing it at $x_{0}=2.0$ as we did previously. For this experiment, we picked $n_{0}=1,000$ and $n_{1}=100$. We additionally trained the conditional density estimator using only $n=100$ data from the high fidelity simulator. All the other settings remain the same.
Table 2: Mean and standard deviation of NLPD and KLD.
ML-NPE TL-NPE NPE (high only)
NLPD $\downarrow$ -0.18 -0.89 -1.21
(0.04) (0.52) (0.53)
KLD $\downarrow$ 0.98 1.24 1.05
(0.08) (0.40) (0.14)
We observe that although our method shows a slight improvement in KLD, it notably underperforms in NLPD compared to using only high-fidelity data. This may be due to our training objective being more unstable and therefore more sensitive to large differences in simulator outputs introduced by the additional parameter $x_{0}$.
B.4 Toggle switch model
@@
We follow the implementation by Key et al. [2025]. Given parameters $\theta=[\alpha_{1},\alpha_{2},\beta_{1},\beta_{2},\mu,\sigma,\gamma]^{\top}$, we can sample from the simulator by $x\sim\widetilde{\mathcal{N}}(\mu+u_{T},\mu\sigma/u_{T}^{\gamma})$, where $\widetilde{\mathcal{N}}$ denotes the truncated normal distribution on $\mathbb{R}_{+}$. Here, $u_{T}$ is given by the discretised-time equation as follows:
@@
$C_{\text{ML-NLE}}$ $=c\left(n_{0}T_{0}+\sum_{l=1}^{L}n_{l}(T_{l}+T_{l-1})\right)$
$C_{\text{NLE}}$ $=cTn.$
This yields a total data generation cost for ML-NLE with $(n_{0},n_{1},n_{2})=(10^{4},500,100)$, and $(T_{0},T_{1},T_{2})=(50,80,300)$ as $C_{\text{MLMC}}=603000$, assuming a unit cost $c=1$. We then allocate the same computational budget to NLE at different fidelities. This results in the following sample sizes: $n_{2}=C_{\text{ML-NLE}}/T_{2}=2010$, $n_{1}=C_{\text{ML-NLE}}/T_{1}=7537$, and $n_{0}=C_{\text{ML-NLE}}/T_{0}=12060$.
Note that $\mathcal{U}$ varies with the fidelity level of the simulator, as it is given by $\mathbb{R}^{2T+1}$. This discrepancy can complicate seed-matching across fidelities. To address this, we define a unified domain for the noise $\mathcal{U}=\bigcup_{l=0}^{L}\mathcal{U}^{l}$ shared across all fidelity levels, and assume that for lower fidelities (smaller $l$), certain components of the input are simply disregarded. For seed-matching, we share $\theta$ and $u_{l}\cap u_{l-1}=u_{l-1}$ where $u_{l}\in\mathcal{U}^{l}\subseteq\mathcal{U},\forall l$.
@@
We use Gaussian mixture density network [Bishop, 1994] with two mixture components. We use MLP with $2$ hidden layers and $20$ units to estimate the parameters: means, variances, and mixture weights, trained for $10,000$ epochs.
@@
Figure 8: Additional results for the Toggle-switch experiment. (a) Histograms of $1000$ seed-matched samples from the high-, the medium-, and the low-fidelity simulators. Note that the difference between the low- and the medium-fidelity simulator is larger than the difference between the medium- and the high-fidelity simulator. This suggests that we should take $n_{1}>n_{2}$. (b) Approximated density and samples from high-fidelity simulator across four simulations. ML-NLE demonstrates superior performance compared to the NLE baselines trained exclusively on high-, medium-, and low-fidelity data with same total simulation budget. In particular, the density estimated by ML-NLE aligns more closely with high-fidelity simulations, highlighting the improved accuracy of our method.
B.4.1 Experiment with different allocation of samples $n_{l}$
We now repeat the same experiment, but with different allocations of $n_{0},n_{1},n_{2}$, keeping the total computational cost roughly the same. Apart from the choice of $n_{0}=10,000,n_{1}=500,n_{2}=100$ that we use in Section˜5.3 (option A) which is guided by our theory, we include two other alternatives: (B) $n_{0}=9260,n_{1}=200,n_{2}=300$ and (C) $n_{0},n_{1},n_{2}=1077$. Option B places more budget on learning the difference between the medium- and the high-fidelity simulator (opposite of option A), and option C allocates equal number of samples for all the terms. The results in Table˜3 indicate that option A is the best performing, indicating performance can be improved by careful assignment of the computational budget led by insights from our theory.
Table 3: Mean and standard deviation of MMD. The results for option A, only high, medium, and low remains the same as in Section˜5.3.
Option A Option B Option C only high only medium only low
MMD ($\downarrow$) 0.16 0.33 0.29 0.43 0.37 0.59
(0.23) (0.26) (0.28) (0.29) (0.45) (0.65)
B.5 Cosmological simulations